# ICONS: Influence Consensus for Vision-Language Data Selection

### v2

@@ B.1 Consistency Analysis
Figure 8: Rel. (%) Across Runs. We show the performance across three different runs for random selection and our ICONS. Our ICONS achieves 98.6$\pm$1.2while the random baseline reaches 95.8$\pm$2.7%.
B.2 Computational Complexity
Computing gradient-based influence requires significant computational resources. In the specialist stage, the complexity scales with both the dataset size $|\mathcal{D}|$ and the gradient dimension $d$. This stage consists of three steps. First, the warm-up training has a complexity of $\mathcal{O}(|\mathcal{D}_{\text{warmup}}|)$. Second, the gradient computation stage has a computational complexity of $\mathcal{O}(|\mathcal{D}|+|\mathcal{D}_{\text{val}}|)$ for forward and backward passes, with storage requirements of $\mathcal{O}(|\mathcal{D}|\cdot d+|\mathcal{D}_{\text{val}}|\cdot d)$ for the gradients. Third (and finally), the influence matrix computation requires $\mathcal{O}(|\mathcal{D}|\cdot|\mathcal{D}_{\text{val}}|\cdot d^{\prime})$ compute cost, where $d^{\prime}$ is the reduced dimension after projection. The generalist stage, focusing on influence consensus across tasks, has lower computational requirements. It begins with threshold computation, requiring $\mathcal{O}(K\cdot|\mathcal{D}|\log|\mathcal{D}|)$ operations for sorting across $K$ tasks. The voting process then takes $\mathcal{O}(K\cdot|\mathcal{D}|)$ compute, followed by a final selection step with complexity $\mathcal{O}(|\mathcal{D}|\log|\mathcal{D}|)$ for sorting the aggregated votes. Storage requirements for this stage are minimal, primarily needing space for the final selected subset.
B.3 Visual Dependency Influence Ranking
Recent work [39] has shown that vision-language tasks vary in their reliance on visual information: task like MMBench [47] depends heavily on visual grounding, while others like SQA-I [25] can be handled primarily through language, showing only a 5% drop in performance when visual input is removed [39]. To take visual dependency of training data into consideration, we further explored gradient-based Visual Dependency Score (VDS). For each data point, we calculate the gradient of the model’s auto-regressive cross-entropy loss with both the original image and a Gaussian noise image $I_{\text{noise}}\sim\mathcal{N}(0,1)$, keeping the text input constant. This quantifies how much the visual component contributes to the model performance. We construct an adapted influence matrix: visual influence matrix $\mathcal{I}_{\texttt{VDS}}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}|}$, which quantifies the visual influence of each training sample ${\bm{z}}_{i}$ on each validation sample ${\bm{z}}^{\prime}_{j}$ with respect to the model’s gradient alignment and visual dependency. $\mathcal{I}_{\text{VDS}}$ is computed as:
Our empirical results demonstrate that VDS-based data selection has varying effectiveness across different vision-language tasks (Tab. 8). The approach shows substantial improvements on tasks requiring strong visual understanding, such as open-ended generation (LLaVA-W Bench: +2.72%) and multiple-choice understanding (MMBench-EN: +1.90%, MMBench-CN: +0.90%). However, tasks that primarily rely on textual reasoning show decreased performance, including SQA-I (-1.84%) and TextVQA (-1.44%). These results align with and extend the findings in Cambrian [39], demonstrating that the effectiveness of VDS corresponds to a task’s visual dependency - tasks that maintain performance without visual inputs show limited or negative impact from VDS-based selection, while visually-dependent tasks benefit significantly. This pattern suggests that VDS effectively identifies training samples where visual information plays an important role in model learning.
Task w/o VDS w/ VDS Delta (%)
VQAv2 [9] 76.3 75.8 -0.66
GQA [13] 60.7 60.9 +0.33
VizWiz [10] 50.1 50.3 +0.40
SQA-I [25] 70.8 69.5 -1.84
TextVQA [37] 55.6 54.8 -1.44
POPE [21] 87.5 86.8 -0.80
MME [7] 1485.7 1489.3 +0.24
MMBench (en) [47] 63.1 64.3 +1.90
MMBench (cn) [47] 55.8 56.3 +0.90
LLaVA-W Bench [23] 66.1 67.9 +2.72
Table 8: Impact of Visual Dependency Score (VDS) on Selection Performance. Performance comparison between influence calculation with and without VDS. VDS shows strongest improvements on LLaVA-W Bench (+2.72%), MMBench (en) (+1.90%), and MMBench (cn) (+0.90%), while tasks like SQA-I (-1.84%), TextVQA (-1.44%), and POPE (-0.80%) show significant performance drops.
Appendix C Algorithm Details
@@
0:  Training dataset $D=\{(x_{i},I_{i},y_{i})\}_{i=1}^{N}$, target tasks $T=\{T_{1},...,T_{K}\}$
0:  Warm-up ratio $r$ (default 5%)
0:  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
1:  for each task $T_{k}\in T$ do
2:     // Step 1: Warm-up Training
3:     Sample warm-up set $D_{\text{warmup}}\subset D$ of size $r|D|$
4:     $f_{\text{warmup}}\leftarrow\text{LoRA}(f_{\text{base}},D_{\text{warmup}})$
5:     // Step 2: Gradient Computation
6:     for each training data $z_{i}\in D$ do
7:        $g_{i}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z_{i}),y_{i})$
8:        $\tilde{g}_{i}\leftarrow\text{Normalize}(Rg_{i})$Random projection}
9:     end for
10:     for each validation data $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
11:        $g^{\prime}_{j}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z^{\prime}_{j}),y^{\prime}_{j})$
12:        $\tilde{g}^{\prime}_{j}\leftarrow\text{Normalize}(Rg^{\prime}_{j})$
13:     end for
14:     // Step 3: Influence Matrix Computation
15:     for each $z_{i}\in D$, $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
16:        $I_{ij}^{k}\leftarrow\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$
17:     end for
@@
20:  end for
21:  return  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
@@
0:  Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
0:  Selection ratio $p$, number of tasks $K$
0:  Selected subset $S\subset D$ of size $m\ll N$
1:  // Compute voting thresholds
2:  for each task $T_{k}\in T$ do
3:     $\tau_{k}\leftarrow(1-p)\text{-th percentile of }\{\bar{I}_{k}(z_{i})\}_{i=1}^{N}$
4:  end for
5:  // Voting process
6:  for each training sample $z_{i}\in D$ do
7:     $I_{\text{vote}}(z_{i})\leftarrow 0$
8:     for each task $T_{k}\in T$ do
9:        $\text{vote}_{k}(z_{i})\leftarrow\mathbbm{1}[\bar{I}_{k}(z_{i})\geq\tau_{k}]$
10:        $I_{\text{vote}}(z_{i})\leftarrow I_{\text{vote}}(z_{i})+\text{vote}_{k}(z_{i})$
11:     end for
12:  end for
@@
Appendix D Future Work
@@
Appendix E Visualizations
@@
Figure 9: VQAv2. Top-left: A sample from VQAv2 [9]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 10: GQA. Top-left: A sample from GQA [13]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 11: VizWiz. Top-left: A sample from VizWiz [10]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 12: SQA. Top-left: A sample from SQA [25]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 13: TextVQA. Top-left: A sample from TextVQA [37]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 14: Pope. Top-left: A sample from Pope [21]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 15: MME. Top-left: A sample from MME [7]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 16: MMBench (en). Top-left: A sample from MMBench (en) [47]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 17: MMBench (cn). Top-left: A sample from MMBench (cn) [47]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 18: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench [23]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 19: Generalist. We show top three influential samples selected after the generalist step.

### v3

@@ D.3 Cross-architecture generalization
Figure 7: Rel. (%) Across Runs. We show the performance across three different runs for random selection and our ICONS.
D.4 Consistency analysis
To evaluate the consistency of ICONS, we conduct three independent runs of the experiments. As shown in Fig. 7, our method demonstrates high consistency across different runs, achieving 98.6$\pm$1.2Rel., which shows a notable improvement over random baseline, which achieves 95.8$\pm$2.7%. The lower standard deviation (1.2% vs 2.7%) further indicates that our approach produces more stable and reliable outcomes compared to the random baseline.
Appendix E Algorithm details
@@
1:Training dataset $D=\{(x_{i},I_{i},y_{i})\}_{i=1}^{N}$, target tasks $T=\{T_{1},...,T_{K}\}$
2:Warm-up ratio $r$ (default 5%)
3:Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
4:for each task $T_{k}\in T$ do
5:   // Step 1: Warm-up Training
6:   Sample warm-up set $D_{\text{warmup}}\subset D$ of size $r|D|$
7:   $f_{\text{warmup}}\leftarrow\text{LoRA}(f_{\text{base}},D_{\text{warmup}})$
8:   // Step 2: Gradient Computation
9:   for each training data $z_{i}\in D$ do
10:      $g_{i}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z_{i}),y_{i})$
11:      $\tilde{g}_{i}\leftarrow\text{Normalize}(Rg_{i})$Random projection}
12:   for each validation data $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
13:      $g^{\prime}_{j}\leftarrow\nabla_{\theta_{w}}L(f_{\text{warmup}}(z^{\prime}_{j}),y^{\prime}_{j})$
14:      $\tilde{g}^{\prime}_{j}\leftarrow\text{Normalize}(Rg^{\prime}_{j})$
15:   // Step 3: Influence Matrix Computation
16:   for each $z_{i}\in D$, $z^{\prime}_{j}\in D_{\text{val}}^{k}$ do
17:      $I_{ij}^{k}\leftarrow\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$
@@
20:return Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
@@
1:Task-specific influence scores $\{\bar{I}_{k}\}_{k=1}^{K}$
2:Selection ratio $p$, number of tasks $K$
3:Selected subset $S\subset D$ of size $m\ll N$
4:// Compute voting thresholds
5:for each task $T_{k}\in T$ do
6:   $\tau_{k}\leftarrow(1-p)\text{-th percentile of }\{\bar{I}_{k}(z_{i})\}_{i=1}^{N}$
7:// Voting process
8:for each training sample $z_{i}\in D$ do
9:   $I_{\text{vote}}(z_{i})\leftarrow 0$
10:   for each task $T_{k}\in T$ do
11:      $\text{vote}_{k}(z_{i})\leftarrow\mathbbm{1}[\bar{I}_{k}(z_{i})\geq\tau_{k}]$
12:      $I_{\text{vote}}(z_{i})\leftarrow I_{\text{vote}}(z_{i})+\text{vote}_{k}(z_{i})$
@@
Appendix F Limitations
Our approach primarily faces one practical limitation: computing gradients for large training datasets is computationally expensive (Appendix §A). This computational overhead could potentially constrain the method’s applicability when working with extremely large-scale datasets. To support broader research community, we release LLaVA-ICONS-133K dataset to help research iteration and model development under resource-constrained settings.
Appendix G Broader impact
Our exploration focuses on scientific understanding and practical applications of vision-language data selection. While our work does not directly imply negative impacts, it may indirectly propagate existing biases present in the original datasets. Therefore, it is important to incorporate rigorous bias-mitigation measurements for data selection. On the positive side, our method enables more efficient and sustainable model development by reducing data redundancy, computational cost while maintaining or even improving performance. Discussion on these critical aspects should remain a priority as we further explore the potential of vision-language data selection.
Appendix H Future work
@@
Appendix I Visualizations
I.1 Representation-based vs. Gradient-based data selection
While ICONS leverages gradient-based influence signals, we explore how representation-based data selection (RDS) performs in the same setting (§3.2). We analyze the top-ranked training examples selected by RDS after the generalist stage in Fig. 8 vs. samples selected by ICONS in Fig. 19. Interestingly, we observe that the representation-based variants often favor training examples with repeated images or instructions, which may dominate the learned representations without contributing to better generalization. Some of the highest-scoring samples under representation-based similarity are duplicated image-question pairs with only the answer choices shuffled. We hypothesize that this is a side effect of the way multimodal representations are constructed—where visually dominant or textually redundant samples occupy high-density regions in embedding space. However, these samples do not necessarily translate into broader utility across tasks, as seen in the performance gap in Tab. 1. This discrepancy raises broader questions about what it means for multimodal data to be diverse. While we leave these questions open for future exploration, our results suggest that gradient-based influence, though computationally more expensive, is better aligned with generalization and multi-task data mixture settings.
Figure 8: Top three samples selected by RDS. These highly-ranked samples are selected due to representation-based similarity but do not necessarily contribute to better generalization, highlighting a key limitation of representation-based selection in multimodal settings.
I.2 Specialists & Generalist
@@
Figure 9: VQAv2. Top-left: A sample from VQAv2 vqav2 . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 10: GQA. Top-left: A sample from GQA gqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 11: VizWiz. Top-left: A sample from VizWiz vizwiz . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 12: SQA. Top-left: A sample from SQA sqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 13: TextVQA. Top-left: A sample from TextVQA textvqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 14: Pope. Top-left: A sample from Pope pope . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 15: MME. Top-left: A sample from MME mme . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 16: MMBench (en). Top-left: A sample from MMBench (en) mmbench . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 17: MMBench (cn). Top-left: A sample from MMBench (cn) mmbench . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 18: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench llava . Remaining panels show top three influential samples selected using the specialist influence ranking step.
Figure 19: Generalist. We show top three influential samples selected after the generalist stage.